CakeCo, Inc. has three operating departments. Information about these departments is listed below. Maintenance is service department at CakeCo that incurred $12,000 of costs during the period. If allocated maintenance cost is based on floor space occupied by each of the operating departments, compute the amount of maintenance cost allocated to the Baking Department.

Mixing Baking Packaging
Direct costs $ 21,000 $ 15,000 $ 9,000
Sq. ft. of space 1,000 1,500 500

Multiple Choice

  • $1,200.

  • $6,000.

  • $4,000.

  • $400.

  • $7,500.
Answer #1

Answer : Allocation of Maintenance to Baking Department = $6,000

Total Maintenance cost = $12,000

Cost Driver = Floor Space

Total Floor Space = (1000+1500+500) = 3000 Sq feet

Allocation of Maintenance cost to Baking Depratment

= (Maintenance cost / Total Floor Space ) * Baking Department Floor Sapce =

=( 12,000/ 3000) * 1,500 = $6,000 (Answer)